OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (anonymous):

equation of the line y=mx+c where m is slope find c value by substituting (–1, 4) in x,y place

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

An alternate method is to use the point-slope form, then convert it to standard form.\[y-4=-(x+1) \implies y=-x+3 \implies x+y=3\]The third equation is in general form, called standard form in some texts.

OpenStudy (anonymous):

Thank you